A rhinoceros is at the origin of coordinates at t1 = 0. For the time interval from 0 to 12 sec, the rhino’s average velocity has x-component -3.8 m/s and y-component 4.9 m/s. At t2 = 12 sec, (a) what are the x- and y-component of the rhino’s position? (b) How far is the rhino from the origin?


1
Expert's answer
2021-07-19T09:50:14-0400

Given:

"v_x=-3.8\:\rm m/s,\quad v_y=4.9\:\rm m/s"

(a)

"{x=v_xt=-3.8*12=-45.6\:\rm m} \\
y=v_yt=4.9.8*12=58.8\:\rm m"

(b)

"d=\sqrt{x^2+y^2}=\sqrt{(-45.6)^2+58.8^2}=74.4\:\rm m"